the sum of the speeds of two trains is 720.7 miles per hour. if the speed of the first train is 7.3 mph faster than the second train, find the speeds of each
Found 2 solutions by ewatrrr, lwsshak3:
Answer by ewatrrr(24785)

(r+7.3) + r = 720.7
2r = 713.4
r = 356.7mph, speed of second train
first train: 364mph

Answer by lwsshak3(11628)

the sum of the speeds of two trains is 720.7 miles per hour. if the speed of the first train is 7.3 mph faster than the second train, find the speeds of each
***
let x=speed of 2nd train
x+7.3=speed of 1st train
x+(x+7.3)=720.7
2x+7.3=720.7
2x=720.7-7.3
2x=713.4
x=356.7
x+7.3=364
speed of 2nd train=356.7 mph
speed of 1st train=364 mph